Grade 12th passMechanics

what is the sum in unit-vector notation of two vectors a=5i+3j & b=-3i+2j? what are the magnitude & direction of a+b ?

Muhammad Mansoor Sarfraz
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear Student
 
Sum = a+b = 2i + 5 j
hence magnitude = (2^2 + 5^2)1/2
 = 291/2
for direction
Theta = tan –1(5/2)
 
regards
Arun (askIITians forum expert)